A) what is the density of Kriptin gas at standard temperature and pessure? The atomic mass of Kripton is 83.80 atomic mass unit. --> 3.74 kg/m^3

B) What is the density of Nitrogen gas under the same conditions?
Answer #1

A)

Assuming ideal gas behavior:
p∙V = n∙R∙T
<=>
V = n∙R∙T/p
So the density is
ρ = m/V = m∙p/(n∙R∙T)
with molar mass
M = m/n
follows
ρ = M∙p / (R∙T)

For Krypton
M = 83.80 g∙mol⁻¹ = 8.38×10⁻² kg∙mol⁻¹
Standard conditions
P = 101325 Pa
T = 273.15 K
Hence,
ρ = 8.38×10⁻² kg∙mol⁻¹ ∙ 101325 Pa / (8.314472 Pa∙m³∙K⁻¹∙mol⁻¹ ∙ 273.15 K)
= 3.74 kg∙mol⁻³

B)

As shown above the density of an ideal gas at same conditions is proportional to its molar mass.
ρ ~ M
=>
ρ(N) / ρ(Kr) = M(N) / M(Kr)
=>
ρ(N) = ρ(Kr) ∙ (M(N)/M(Kr))
The molar mass of methane is:
M(N) = 14.00 g∙mol⁻¹
=>
ρ(N) = 3.74 kg∙mol⁻³ ∙ (14 g∙mol⁻¹ / 83.8 g∙mol⁻¹)
= 0.6248 kg∙mol⁻³